OpenStudy (anonymous):

how would you write a polynomial with rational coefficients having roots 3, 3+i, and 3-i?

OpenStudy (dumbcow):

x = 3 +- i x-3 = i (x-3)^2 = -1 x^2 -6x +10 = 0 polynomial --> (x-3)(x^2 -6x +10)